OpenStudy (anonymous):

what's the domain of \[x^3y^2z \sqrt{10-x-y-z}\]

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (atlas):

domain within real numbers??

OpenStudy (anonymous):

yes sir

OpenStudy (atlas):

So to keep the given expression within real numbers - the expression within the square root cannot be negative........What is the condition for this?

OpenStudy (anonymous):

10-x-y-z >= 0

OpenStudy (atlas):

right or, x+y+z<=0

OpenStudy (anonymous):

ok what's next..?

OpenStudy (anonymous):

isn't it x+y+z <= 10 ?

OpenStudy (atlas):

yeah i m sry abt that

OpenStudy (atlas):

Now you have to find the maximum and minimum values of the given expression within the given constraint

OpenStudy (atlas):

For finding maximum value x and z have to be positive while y can be negative...........I hope you can guess why?

OpenStudy (anonymous):

.